The 200 m free style women’s swimming gold medal at Seol Olympic 1988 went to Heike Friendrich of East Germany when she set a ne Olympic record of 1 minute and 57.56 seconds. Assume that she covered most of the distance with a uniform speed and had to exert 460 W to maintain her speed. Calculate the average force of resistance offered by the water during the swim.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. t = 1 min 57.56 sec = 11.56 sec, p= 400 W, s =200 m p = W/t, Work w = pt = 460 * 117.56 J Again, W = FS = 460 * 117.56/200 = 270.3 N = 170 N